OpenStudy (anonymous):

Help Please !!! Find h′(2) given that f(2) = −1, f′(2) = 4, g(2) = 1, and g′(2) = −5. solve each one a. h(x) = 5f(x) + 2g(x). b. h(x) = f(x)g(x). c. h(x) = f(x)/g(x). d. h(x) = ln f(x).

OpenStudy (anonymous):

please help someone

OpenStudy (ash2326):

@hooverst We'll solve each option a)h(x)=5f(x)+2g(x) Let's differentiate this \[h'(x)=5f'(x)+2g'(x)\] we want to find h'(2) so put x=2 \[h'(2)=5f'(2)+2g'(2)\] Can you solve now?

OpenStudy (anonymous):

@ash2326 good :)

OpenStudy (anonymous):

yes

OpenStudy (jiteshmeghwal9):

plug x as 2 in given expression as\[h(2)=5f(2)+2g(2)\]\[h(2)=5(-1)+2(1)\]\[h(2)=-5+2=-3\]

OpenStudy (ash2326):

Can you differentiate b? \[f(x)=h(x)\times g(x)\]

OpenStudy (jiteshmeghwal9):

&\[h(2)=5(2)+2(-5)\]\[h(2)=10+(-10)=0\]

OpenStudy (anonymous):

ok got it.

OpenStudy (ash2326):

@hooverst Did you try the third option?

OpenStudy (anonymous):

I'm working on it now I would divide right?

OpenStudy (anonymous):

the one above how is that 5(2) instead of 4(2)

OpenStudy (ash2326):

\[h'(x)=f'(x)\times g(x)+f(x)\times g'(x)\] x=2 \[h'(2)=f'(2)\times g(2)+f(2)\times g'(2)\] \[h'(2)=4\times 1+(-1\times -5)=4+5\]

OpenStudy (ash2326):

@hooverst do you know the quotient rule of division?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would the anserw be 0 for c ?

OpenStudy (anonymous):

I'm not sure if I did it right.